A hydraulic press has a small piston of diameter =3.4 cm and the large piston of diameter=30 cm. If the force applied to the small piston (F1) is 395 N.

What is the mass of the car that can be raised at the large piston (in kg)?


1
Expert's answer
2021-11-25T10:14:21-0500
"\frac{mg}{R^2}=\frac{F}{r^2}\\\frac{9.8m}{30^2}=\frac{395}{3.4^2}\\m=3138\ kg"